# Quantum Dynamic Programming

### v1

@@ II.2 Hermitian-preserving map exponentiation and its extension
The property of the swap operator $\hat{S}^{2}=1$ facilitates the derivation of an explicit form of Eq. (39):
^s$\hat{\mathcal{E}}^{(\rho)}_{s}\left(\sigma\right)=\cos^{2}(s)\sigma-i\sin(s)\cos(s)[\rho,\sigma]+\sin^{2}(s)\rho\ .$ (40)
Since Def. 6 is a special case of Def. 5, it approximates the corresponding unitary channel $\hat{\mathbf{E}}^{(\operatorname{id},\rho)}_{s}$ according to Eq. (32), which we denote for simplicity as $\hat{\mathbf{E}}^{(\rho)}_{s}(\sigma)\coloneq e^{-i\rho s}\sigma e^{i\rho s}$. As a special instance of Eq. (37) when $\|\hat{N}\|_{\infty}=\|\hat{S}\|_{\infty}=1$,
$\frac{1}{2}\left\lVert\left(\hat{\mathcal{E}}^{(\rho)}_{s/M}\right)^{M}-\hat{\mathbf{E}}^{(\rho)}_{s}\right\rVert_{\operatorname{Tr}}\leq\frac{1}{2}\left\lVert\left(\hat{\mathcal{E}}^{(\rho)}_{s/M}\right)^{M}-\hat{\mathbf{E}}^{(\rho)}_{s}\right\rVert_{\diamond}=\mathcal{O}\left(\frac{s^{2}}{M}\right)\ .$ (41)
@@
Using Lemma 4, a linear map $g$, such that $g(\rho^{\otimes d})=f(\rho)$ for any $f$, exists. Then we can approximate $e^{ig(\rho)}$ using Hermitian-preserving map exponentiations ^s$\hat{\mathcal{E}}^{(\hat{G},\rho^{\otimes d})}_{s}$, defined in Def. 5, where $\hat{G}$ is the partial transpose of the Choi matrix of $g$. Following Eq. (38), we take $s=M^{-1}$ and concatenate the channel ^$\hat{\mathcal{E}}^{(\hat{G},\rho^{\otimes d})}_{M^{-1}}$ for $M$ times, which leads to the implementation error $\mathcal{O}(\lVert\hat{G}\rVert_{\infty}^{2}M^{-1})$. Setting $M=\mathcal{O}(\lVert\hat{G}\rVert_{\infty}^{2}\epsilon^{-1})$, we can achieve the approximation with arbitrarily small error $\epsilon$. Each channel ^s$\hat{\mathcal{E}}^{(\hat{G},\rho^{\otimes d})}_{s}$ requires a circuit of depth $\mathcal{O}(1)$ and $\mathcal{O}(d)$ copies of $\rho$. Hence, $M$ concatenations of such channels cost $\mathcal{O}(d\epsilon^{-1}\lVert\hat{G}\rVert_{\infty}^{2})$ copies of $\rho$ and the circuit depth $\mathcal{O}(\epsilon^{-1}\lVert\hat{G}\rVert_{\infty}^{2})$. ∎
@@ III Fast spectral convergence and efficient quantum dynamic programming
We begin by illustrating a class of fixed-point iterations that admit efficient QDP implementation. Our recursions are unitary; hence, it is natural to consider iterations with a fixed-point shared by isospectral states (states sharing the same eigenvalue spectrum). In this section, we denote a generic quantum recursion in the form Eq. (27) as $\hat{U}^{(\rho)}$.
@@
Consider a quantum recursion $\hat{U}^{(\rho)}$ that defines a fixed-point iteration
$\rho\mapsto\hat{U}^{(\rho)}\rho\left(\hat{U}^{(\rho)}\right)^{\dagger}\ .$ (42)
@@
$\delta_{n+1}\leq h(\delta_{n})<\delta_{n},$ (43)
@@
If a quantum recursion $\hat{U}^{(\rho)}$ satisfies fast spectral convergence for some initial state $\rho_{0}$, then the QDP implementation of $N$ iterations prepares a state $\rho^{\prime}_{N}$, whose distance to the fixed-point $\delta_{\rho^{\prime}_{N}}\leq h^{N}(\delta_{0})+\epsilon$, by using a circuit with depth $\mathcal{O}(N^{2}\epsilon^{-1})$.
@@
Let us denote the unitary channel corresponding to the operator $\hat{U}^{(\rho)}$ as $\hat{\mathbf{U}}^{(\rho)}$. The recursion is spectrally converging. Hence, there exists a fixed-point $\tau$, such that the distance $\frac{1}{2}\lVert\tau-\rho\rVert_{1}\eqcolon\delta_{\rho}$ follows
$\delta_{\hat{\mathbf{U}}^{(\rho)}(\rho)}=\frac{1}{2}\left\lVert\tau-\hat{\mathbf{U}}^{(\rho)}(\rho)\right\rVert_{1}\leq h(\delta_{\rho})<\delta_{\rho}\ .$ (44)
@@
$h(\delta+\epsilon)\leq h(\delta)+r\epsilon\ ,$ (45)
@@
From Lemma 7, we are able to locally accurately implement the unitary channel $\hat{\mathbf{U}}^{(\rho)}$ by a non-unitary channel $\hat{\mathcal{U}}^{(\rho)}$, such that
$\frac{1}{2}\left\lVert\hat{\mathbf{U}}^{(\rho)}-\hat{\mathcal{U}}^{(\rho)}\right\rVert_{\operatorname{Tr}}\leq\eta,$ (46)
@@
$\rho^{\prime}_{n}=\hat{\mathcal{U}}^{(\rho^{\prime}_{n-1})}\left(\rho^{\prime}_{n-1}\right)\ .$ (47)
@@
We additionally consider a sequence of states $\{\tilde{\rho}_{n}\}_{n}$, defined by the exact unitary recursion with an erroneous instruction state $\tilde{\rho}_{n}=\hat{\mathbf{U}}^{(\rho^{\prime}_{n-1})}(\tilde{\rho}_{n-1})$ starting from $\tilde{\rho}_{0}=\rho_{0}$. Hence, $\mathrm{spec}\{\tilde{\rho}_{n}\}=\mathrm{spec}\{\tau\}$. Let us denote $\frac{1}{2}\lVert\tilde{\rho}_{n}-\rho^{\prime}_{n}\rVert_{1}\eqcolon\varepsilon_{n}$ with $\varepsilon_{0}=0$. We first analyze how $\varepsilon_{n}$ scales. Observe that
$\varepsilon_{n+1}=\frac{1}{2}\left\lVert\hat{\mathbf{U}}^{(\rho^{\prime}_{n})}(\tilde{\rho}_{n})-\rho^{\prime}_{n+1}\right\rVert_{1}\leq\frac{1}{2}\left\lVert\hat{\mathbf{U}}^{(\rho^{\prime}_{n})}(\tilde{\rho}_{n})-\hat{\mathbf{U}}^{(\rho^{\prime}_{n})}(\rho^{\prime}_{n})\right\rVert_{1}+\frac{1}{2}\left\lVert\hat{\mathbf{U}}^{(\rho^{\prime}_{n})}(\rho^{\prime}_{n})-\hat{\mathcal{U}}^{(\rho^{\prime}_{n})}(\rho^{\prime}_{n})\right\rVert_{1}\leq\varepsilon_{n}+\eta\ ,$ (48)
from the triangle inequality, the unitary invariance of the trace norm, and Eq. (46). Therefore, $\varepsilon_{n}$ scales linearly as $\varepsilon_{n}\leq n\eta$. The triangle inequality between $\tau$, $\rho^{\prime}_{n}$, and $\rho_{n}$ yields $\delta_{\rho^{\prime}_{n}}\leq\delta_{\tilde{\rho}_{n}}+\varepsilon_{n}$, which connects the quantity of our interest $\delta_{\rho^{\prime}_{n}}$ and the term $\delta_{\tilde{\rho}_{n}}$ derived from a state $\tilde{\rho}_{n}$ isospectral to the desired fixed-point $\tau$.
@@
$\left\lVert\hat{\mathbf{U}}^{(\rho)}-\hat{\mathbf{U}}^{(\sigma)}\right\rVert_{\operatorname{Tr}}\leq C\lVert\rho-\sigma\rVert_{1}$ (49)
@@
$\delta_{\tilde{\rho}_{n+1}}\leq\frac{1}{2}\left\lVert\tau-\hat{\mathbf{U}}^{(\tilde{\rho}_{n})}\left(\tilde{\rho}_{n}\right)\right\rVert_{1}+\frac{1}{2}\left\lVert\hat{\mathbf{U}}^{(\tilde{\rho}_{n})}\left(\tilde{\rho}_{n}\right)-\hat{\mathbf{U}}^{(\rho^{\prime}_{n})}\left(\tilde{\rho}_{n}\right)\right\rVert_{1}\leq h\left(\delta_{\tilde{\rho}_{n}}\right)+\frac{1}{2}\left\lVert\hat{\mathbf{U}}^{(\tilde{\rho}_{n})}-\hat{\mathbf{U}}^{(\rho^{\prime}_{n})}\right\rVert_{\operatorname{Tr}}\leq h\left(\delta_{\tilde{\rho}_{n}}\right)+C\varepsilon_{n},$ (50)
where we use the triangle inequality for the first, the assumption Eq. (44) for the second, and Eq. (49) for the last inequality. The property of $h$, Eq. (45) gives
$\delta_{\tilde{\rho}_{n}}\leq h\left(\delta_{\tilde{\rho}_{n-1}}\right)+C\varepsilon_{n-1}\leq h\left(h\left(\delta_{\tilde{\rho}_{n-2}}\right)+C\varepsilon_{n-2}\right)+C\varepsilon_{n-1}\leq h\circ h\Big{(}\delta_{\tilde{\rho}_{n-2}}\Big{)}+rC\varepsilon_{n-2}+C\varepsilon_{n-1}\ ,$ (51)
@@
$\delta_{\tilde{\rho}_{N}}\leq h^{N}\left(\delta_{\rho_{0}}\right)+\sum_{n=0}^{N-1}r^{N-n-1}(Cn\eta)=h^{N}\left(\delta_{\rho_{0}}\right)+\mathcal{O}(N\eta)\ .$ (52)
@@
$\delta_{\rho^{\prime}_{N}}=h^{N}\left(\delta_{\rho_{0}}\right)+\mathcal{O}\big{(}N\eta\big{)}\ .$ (53)

### v2

@@ II.2 Hermitian-preserving map exponentiation and its extension
The property of the swap operator $\hat{S}^{2}=1$ facilitates the derivation of an explicit form of Eq. (S26):
^s$\hat{\mathcal{E}}^{[\rho]}_{s}\left(\sigma\right)=\cos^{2}(s)\sigma-i\sin(s)\cos(s)[\rho,\sigma]+\sin^{2}(s)\rho\ .$ (S27)
Since Def. 6 is a special case of Def. 5, it approximates the corresponding unitary channel $\hat{\mathbf{E}}^{[\operatorname{id},\rho]}_{s}$ according to Eq. (S19), which we denote for simplicity as $\hat{\mathbf{E}}^{[\rho]}_{s}(\sigma)\coloneq e^{-i\rho s}\sigma e^{i\rho s}$. As a special instance of Eq. (S24) when $\|\hat{N}\|_{\infty}=\|\hat{S}\|_{\infty}=1$,
$\frac{1}{2}\left\lVert\left(\hat{\mathcal{E}}^{[\rho]}_{s/M}\right)^{M}-\hat{\mathbf{E}}^{[\rho]}_{s}\right\rVert_{\operatorname{Tr}}\leq\frac{1}{2}\left\lVert\left(\hat{\mathcal{E}}^{[\rho]}_{s/M}\right)^{M}-\hat{\mathbf{E}}^{[\rho]}_{s}\right\rVert_{\diamond}=\mathcal{O}\left(\frac{s^{2}}{M}\right)\ .$ (S28)
@@
Using Lemma 4, a linear map $g$, such that $g(\rho^{\otimes d})=f(\rho)$ for any $f$, exists. Then we can approximate $e^{ig(\rho)}$ using Hermitian-preserving map exponentiations ^s$\hat{\mathcal{E}}^{[\hat{G},\rho^{\otimes d}]}_{s}$, defined in Def. 5, where $\hat{G}$ is the partial transpose of the Choi matrix of $g$. Following Eq. (S25), we take $s=M^{-1}$ and concatenate the channel ^$\hat{\mathcal{E}}^{[\hat{G},\rho^{\otimes d}]}_{M^{-1}}$ for $M$ times, which leads to the implementation error $\mathcal{O}(\lVert\hat{G}\rVert_{\infty}^{2}M^{-1})$. Setting $M=\mathcal{O}(\lVert\hat{G}\rVert_{\infty}^{2}\epsilon^{-1})$, we can achieve the approximation with arbitrarily small error $\epsilon$. Each channel ^s$\hat{\mathcal{E}}^{[\hat{G},\rho^{\otimes d}]}_{s}$ requires a circuit of depth $\mathcal{O}(1)$ and $\mathcal{O}(d)$ copies of $\rho$. Hence, $M$ concatenations of such channels cost $\mathcal{O}(d\epsilon^{-1}\lVert\hat{G}\rVert_{\infty}^{2})$ copies of $\rho$ and the circuit depth $\mathcal{O}(\epsilon^{-1}\lVert\hat{G}\rVert_{\infty}^{2})$. ∎
@@ III Fast spectral convergence and efficient quantum dynamic programming
We begin by illustrating a class of fixed-point iterations that admit efficient QDP implementation. Our recursions operate on general mixed states but unitary; hence, it is natural to consider iterations with a fixed-point shared by isospectral states (states sharing the same eigenvalue spectrum). In this section, we denote a generic quantum recursion in the form Eq. (S14) as $\hat{U}^{[\rho]}$.
@@
Consider a quantum recursion $\hat{U}^{[\rho]}$ that defines a fixed-point iteration
$\rho\mapsto\hat{U}^{[\rho]}\rho\left(\hat{U}^{[\rho]}\right)^{\dagger}\ .$ (S29)
@@
$\delta_{n+1}\leq h(\delta_{n})<\delta_{n},$ (S30)
@@
III.1 Quantum dynamic programming with a mixed initial state
We first establish a theorem that demonstrates the efficiency of quantum dynamic programming when the initial and all the instruction states are mixed.
@@ III.1 Quantum dynamic programming with a mixed initial state
If a quantum recursion $\hat{U}^{[\rho]}$ satisfies fast spectral convergence for some initial state $\rho_{0}$, then the QDP implementation of $N$ iterations prepares a state $\rho^{\prime}_{N}$, whose distance to the fixed-point $\delta_{\rho^{\prime}_{N}}\leq h^{N}(\delta_{0})+\epsilon$, by using a circuit with depth $\mathcal{O}(N^{2}\epsilon^{-1})$.
@@
Let us denote the unitary channel corresponding to the operator $\hat{U}^{[\rho]}$ as $\hat{\mathbf{U}}^{[\rho]}$. The recursion is spectrally converging. Hence, there exists a fixed-point $\tau$, such that the distance $\frac{1}{2}\lVert\tau-\rho\rVert_{1}\eqcolon\delta_{\rho}$ follows
$\delta_{\hat{\mathbf{U}}^{[\rho]}(\rho)}=\frac{1}{2}\left\lVert\tau-\hat{\mathbf{U}}^{[\rho]}(\rho)\right\rVert_{1}\leq h(\delta_{\rho})<\delta_{\rho}\ .$ (S31)
@@
$h(\delta+\epsilon)\leq h(\delta)+r\epsilon\ ,$ (S32)
@@
From Lemma 7, we are able to locally accurately implement the unitary channel $\hat{\mathbf{U}}^{[\rho]}$ by a non-unitary channel $\hat{\mathcal{U}}^{[\rho]}$, such that
$\frac{1}{2}\left\lVert\hat{\mathbf{U}}^{[\rho]}-\hat{\mathcal{U}}^{[\rho]}\right\rVert_{\operatorname{Tr}}\leq\eta,$ (S33)
@@
$\rho^{\prime}_{n}=\hat{\mathcal{U}}^{[\rho^{\prime}_{n-1}]}\left(\rho^{\prime}_{n-1}\right)\ .$ (S34)
@@
We additionally consider a sequence of states $\{\tilde{\rho}_{n}\}_{n}$, defined by the exact unitary recursion with an erroneous instruction state $\tilde{\rho}_{n}=\hat{\mathbf{U}}^{[\rho^{\prime}_{n-1}]}(\tilde{\rho}_{n-1})$ starting from $\tilde{\rho}_{0}=\rho_{0}$. Hence, $\mathrm{spec}\{\tilde{\rho}_{n}\}=\mathrm{spec}\{\tau\}$. Let us denote $\frac{1}{2}\lVert\tilde{\rho}_{n}-\rho^{\prime}_{n}\rVert_{1}\eqcolon\varepsilon_{n}$ with $\varepsilon_{0}=0$. We first analyze how $\varepsilon_{n}$ scales. Observe that
$\varepsilon_{n+1}=\frac{1}{2}\left\lVert\hat{\mathbf{U}}^{[\rho^{\prime}_{n}]}(\tilde{\rho}_{n})-\rho^{\prime}_{n+1}\right\rVert_{1}\leq\frac{1}{2}\left\lVert\hat{\mathbf{U}}^{[\rho^{\prime}_{n}]}(\tilde{\rho}_{n})-\hat{\mathbf{U}}^{[\rho^{\prime}_{n}]}(\rho^{\prime}_{n})\right\rVert_{1}+\frac{1}{2}\left\lVert\hat{\mathbf{U}}^{[\rho^{\prime}_{n}]}(\rho^{\prime}_{n})-\hat{\mathcal{U}}^{[\rho^{\prime}_{n}]}(\rho^{\prime}_{n})\right\rVert_{1}\leq\varepsilon_{n}+\eta\ ,$ (S35)
from the triangle inequality, the unitary invariance of the trace norm, and Eq. (S33). Therefore, $\varepsilon_{n}$ scales linearly as $\varepsilon_{n}\leq n\eta$. The triangle inequality between $\tau$, $\rho^{\prime}_{n}$, and $\rho_{n}$ yields $\delta_{\rho^{\prime}_{n}}\leq\delta_{\tilde{\rho}_{n}}+\varepsilon_{n}$, which connects the quantity of our interest $\delta_{\rho^{\prime}_{n}}$ and the term $\delta_{\tilde{\rho}_{n}}$ derived from a state $\tilde{\rho}_{n}$ isospectral to the desired fixed-point $\tau$.
@@
$\left\lVert\hat{\mathbf{U}}^{[\rho]}-\hat{\mathbf{U}}^{[\sigma]}\right\rVert_{\operatorname{Tr}}\leq C\lVert\rho-\sigma\rVert_{1}$ (S36)
@@
$\delta_{\tilde{\rho}_{n+1}}\leq\frac{1}{2}\left\lVert\tau-\hat{\mathbf{U}}^{[\tilde{\rho}_{n}]}\left(\tilde{\rho}_{n}\right)\right\rVert_{1}+\frac{1}{2}\left\lVert\hat{\mathbf{U}}^{[\tilde{\rho}_{n}]}\left(\tilde{\rho}_{n}\right)-\hat{\mathbf{U}}^{[\rho^{\prime}_{n}]}\left(\tilde{\rho}_{n}\right)\right\rVert_{1}\leq h\left(\delta_{\tilde{\rho}_{n}}\right)+\frac{1}{2}\left\lVert\hat{\mathbf{U}}^{[\tilde{\rho}_{n}]}-\hat{\mathbf{U}}^{[\rho^{\prime}_{n}]}\right\rVert_{\operatorname{Tr}}\leq h\left(\delta_{\tilde{\rho}_{n}}\right)+C\varepsilon_{n},$ (S37)
where we use the triangle inequality for the first, the assumption Eq. (S31) for the second, and Eq. (S36) for the last inequality. The property of $h$, Eq. (S32) gives
$\delta_{\tilde{\rho}_{n}}\leq h\left(\delta_{\tilde{\rho}_{n-1}}\right)+C\varepsilon_{n-1}\leq h\left(h\left(\delta_{\tilde{\rho}_{n-2}}\right)+C\varepsilon_{n-2}\right)+C\varepsilon_{n-1}\leq h\circ h\Big{(}\delta_{\tilde{\rho}_{n-2}}\Big{)}+rC\varepsilon_{n-2}+C\varepsilon_{n-1}\ ,$ (S38)
@@
$\delta_{\tilde{\rho}_{N}}\leq h^{N}\left(\delta_{\rho_{0}}\right)+\sum_{n=0}^{N-1}r^{N-n-1}(Cn\eta)=h^{N}\left(\delta_{\rho_{0}}\right)+\mathcal{O}(N\eta)\ .$ (S39)
@@
$\delta_{\rho^{\prime}_{N}}=h^{N}\left(\delta_{\rho_{0}}\right)+\mathcal{O}\big{(}N\eta\big{)}\ .$ (S40)